# Vortical waves in a quantum fluid with vector, axial and helical charges. II. Dissipative effects

### v1

@@ Abstract
In this paper, we consider the effect of interactions on the local, average polarization of a quantum plasma of massless fermion particles characterized by vector, axial, and helical quantum numbers. Due to the helical and axial vortical effects, perturbations in the vector charge in a rotating plasma can lead to chiral and helical charge transfer along the direction of the vorticity vector. At the same time, interactions between the plasma constituents lead to the dissipation of the helical charge through helicity-violating pair annihilation (HVPA) processes and of the axial charge through the axial anomaly. We will discuss separately a QED-like plasma, in which we ignore background electromagnetic fields and thus the axial charge is roughly conserved, as well as a QCD-like plasma, where instanton effects lead to the violation of the axial charge conservation, even in the absence of background chromomagnetic fields. The non-conservation of helicity and chirality leads to a gapping of the Helical, Axial and mixed, Axial-Helical vortical waves that prevents their infrared modes from propagating. On the other hand, usual dissipative effects, such as charge diffusion, lead to significant damping of ultraviolet modes. We end this paper with a discussion of the regimes where these vortical waves may propagate.
@@ 1 Introduction
Vortical backgrounds – represented, for example, by whirlpools in fluids or plasmas – host a similar class of hydrodynamic phenomenon known under the common name of Chiral Vortical Effects Vilenkin:1978is ; Vilenkin:1979ui . These vortical transport effects generate intertwined oscillations of vector and axial currents, as well as their charges, propagating along the rotation axis Jiang:2015cva . The resulting Chiral Vortical Waves, which are supported by the presence of the chiral anomaly and its mixed chiral-gravitational counterpart Landsteiner:2011cp , were suggested to exist in rotating quark-gluon plasmas that are created in highly-vortical noncentral heavy-ion collisions Jiang:2015cva .
@@
The picture described above traditionally represents a chiral fluid as a two-component system that includes only vector and axial degrees of freedom that propagate coherently. However, at this point, it is worth noticing that fermions possess a third type of quantum number associated with the fermion helicity, which is distinct from the vector and axial charges Pal:2010ih . Helicity and chirality are distinguishable physical properties of fermions that are often inaccurately identified with each other Ambrus:2019ayb ; Ambrus:2019khr . The presence of this additional third characteristic of the fermion ensemble leads to the associated helical vortical transport phenomena that affect, in particular, the spectrum of the hydrodynamic modes Ambrus:2019khr . [^3] Here, for completeness, we mention that similar, helicity-related nontrivial transport and associated hydrodynamic modes in the magnetic field background has been uncovered in Refs. Ambrus:2023erf . In this paper, we concentrate only on rotating systems.
@@
In our companion paper HVWp1 , we have been exploring the role of helical degrees of freedom in vortical chiral fluids in an academic dissipationless limit when both axial and helical quantum numbers are conserved. This study has confirmed the presence of the Helical Vortical Wave Ambrus:2019khr , which corresponds to a gapless hydrodynamic excitation acting mainly in the helical and vector sectors of charge densities. Surprisingly, it also has uncovered yet another gapless wave in the system: the Axial Vortical Wave, which is driven by a purely axial wave accompanied, in the presence of vector (baryon) or helical (spin-polarized) background, by coherent fluctuations in axial and helical/vector charges and their current. At high vector (baryon) densities, where the fluid becomes degenerate, both the Helical and the Axial Vortical Waves merge into a hydrodynamic excitation, the Axial-Helical Vortical Wave. The fascinating feature of the Axial Vortical Wave is its spatial non-reciprocity, which makes this hydrodynamic mode distinguishable from the usual acoustic waves: a pure Axial Vortical Wave propagates only along the direction of the angular velocity while is unable to travel in the opposite direction.
@@
This paper aims to study the fate of the hydrodynamic vortical excitations in realistic plasmas, which are characterized by axial and helical charge relaxation as well as kinetic damping effects. A similar analysis has been extensively applied to determine possible experimentally observable signatures Ke:2012qb ; STAR:2015wza of a similar excitation, the Chiral Magnetic Wave, that emerges in the magnetic field background and incorporates oscillations of the vector and axial charges Newman:2005hd ; Kharzeev:2010gd . In addition to the standard axial charge relaxation effects – related to the fact that the axial charge is not a conserved quantity in quantum field theory – the vector sector is also subjected to relaxation effects caused by the electromagnetic backreaction Kharzeev:2010gd . The latter phenomenon is supported by a finite electrical conductivity in plasma that leads to an efficient screening of the local vector charge fluctuations carrying nonvanishing electric charge density. As a result, the Chiral Magnetic Wave is essentially “overdamped” in realistic environments such as quark-gluon plasma: the relaxation length of the wave appears to be shorter than the half-period of the wave oscillation Shovkovy:2018tks . [^4] A recent work Ahn:2024ozz highlights the regimes, in which the Chiral Magnetic Wave can still be observed in a condensed matter setup as an underdamped collective mode in a Weyl semimetal which hosts relativistic fermionic quasiparticles similarly to quark-gluon plasma. In the vector-chiral sector of rotating plasma, a related analysis has been performed in Ref. Rybalka:2018uzh , where various waves were analyzed, including the longitudinal sound wave, a circularly polarized vortical wave and diffusive modes that involve an oscillating vector charge density.
@@
Here, we include in the analysis the helical charge, which enriches the spectrum of the hydrodynamic excitations. In particular, the Helical Vortical Wave Ambrus:2019khr involves the oscillations of the helical and vector charge densities, and, therefore, our analysis gives a higher, conservative bound of the wave propagation length caused by the helical charge relaxation and kinetic dissipation effects. The electromagnetic backreaction of the vector charge density, not considered in this paper, will restrict this length even further. On the other hand, the Axial and Axial-Helical Vortical waves, found in the companion paper HVWp1 , are driven essentially by electrically neutral helical and axial charge densities, and therefore, they are not subjected to the backreaction coming from the electromagnetic sector associated with the vector charge fluctuations.
@@
One may argue that the helical degree of freedom is as good (bad) as the chiral charge of massive fermions: neither of them is conserved in the physically relevant theories, such as massive QED or QCD. Chirality is not conserved for massless fermions either, due to instanton-like effects that are present in the QCD medium even in the absence of background chromomagnetic fields. In the relaxation time approximation, the pertinence of these quantum numbers for the dynamics of the system is determined by the relative order of magnitudes of the axial, $\tau_{A}$, and helical, $\tau_{H}$, relaxation times, which show how fast these charges dissolve in the system. Similarly to the axial (chiral) charge Ruggieri:2016asg ; Astrakhantsev:2019zkr , the relaxation of the helical charge might also be a rather fast process Ambrus:2020oiw . In addition to the charge relaxation times, the system is also characterized by the kinetic relaxation time $\tau_{R}$, which encodes how fast fluctuations in the thermodynamic characteristics of the system (such as pressure and energy density) relax towards the thermodynamic equilibrium. In general, all three relaxation times, $\tau_{R}$, $\tau_{A}$, and $\tau_{H}$, are independent of each other. An expression for $\tau_{H}$ was derived in Ref. Ambrus:2019khr for a neutral plasma. In this paper, we will extend this derivation for the case of a finite vector imbalance, encoding a difference between the number of particles and anti-particles in the system, which is necessary in order to assess its relevance in the limit of a degenerate (dense) system. A similar calculation for the axial relaxation time is beyond the scope of this work; as such, an analysis must be performed in fully non-perturbative QCD using, e.g., lattice methods.
@@
The structure of the paper is as follows. In Sec. 2, we briefly present the setup for the study of linear perturbations in the quantum plasma with vector/axial/helical degrees of freedom, following the approach in Ref. HVWp1 , which we extend to the case when the helical and/or axial charges are not conserved. Our focus is on the case of an unpolarized background plasma, in which both the helical and axial chemical potentials vanish.
In Sec. 3, we take into account the non-conservation of the helical charge stemming from local processes using a relaxation time approximation. Such a scenario is relevant for a massless QED plasma, where both vector and axial charges are conserved due to an unbroken U(1) gauge symmetry and the vector nature of inter-particle interactions, respectively. The Chiral Vortical Wave is then recovered in the limit where the helical charge dissipates instantaneously. An interplay between Chiral and Helical Vortical Waves, as well as a distinct role played by the Axial Vortical Wave, are discussed.
@@ 2 Hydrodynamic description of waves in V/A/H fluids
In this section, we briefly review the anomalous transport in the presence of rotation. In Subsec. 2.1, we review the constitutive equations for the energy-momentum tensor and the charge currents for a Dirac fluid with vector, axial and helical imbalance, undergoing rigid rotation, as seen in the Landau (or energy) frame. We then summarize the treatment of small perturbations around this rigidly rotating state in Subsec. 2.3.
@@ 2.1 Landau frame decomposition for slow rotation
$T^{\mu\nu}\simeq E_{\beta}u^{\mu}u^{\nu}-P_{\beta}\Delta^{\mu\nu}+u^{\mu}W_{\beta}^{\nu}+u^{\nu}W_{\beta}^{\mu},\quad J^{\mu}_{\ell}\simeq Q_{\ell}u^{\mu}+V_{\beta}^{\mu},$ (1)
where $W_{\beta}^{\mu}=\sigma^{\omega}_{\varepsilon;\beta}\omega^{\mu}$ and $V^{\mu}=\sigma^{\omega}_{\ell;\beta}\omega^{\mu}$ represent the heat flux and diffusion currents in the fluid rest frame, $\omega^{\mu}=\frac{1}{2}\varepsilon^{\mu\alpha\beta\gamma}u_{\alpha}\partial_{\beta}u_{\gamma}\simeq\Omega\delta^{\mu}{}_{z}$ is the vorticity four-vector and we neglected terms of quadratic or higher order in $\Omega$. In the above expression, $u^{\mu}\partial_{\mu}=\Gamma(\partial_{t}+\Omega\partial_{\varphi})$ is the four-velocity of a rigidly-rotating fluid, which defines the so-called $\beta$ hydrodynamic frame. The energy density $E_{\beta}$ is related to the pressure $P$ through the ultrarelativistic equation of state, $E=3P$. The charge densities $Q_{\ell}$ and the heat $\sigma^{\omega}_{\varepsilon;\beta}$ and charge $\sigma^{\omega}_{\ell;\beta}$ vortical conductivities, measured in the $\beta$ frame, can be obtained from the thermodynamic pressure $P$ via
$Q_{\ell}=\frac{\partial P}{\partial\mu_{\ell}},\qquad\sigma^{\omega}_{\varepsilon;\beta}=Q_{A},\qquad\sigma^{\omega}_{\ell}=\frac{1}{2}\frac{\partial Q_{\ell}}{\partial\mu_{A}}.$ (2)
To linear order in $\Omega$, one can employ the Landau hydrodynamic frame, by which the four-velocity $u_{L}^{\mu}$ satisfies $T^{\mu}{}_{\nu}u_{L}^{\nu}=Eu_{L}^{\mu}$, with $E$ being the energy density and $T^{\mu\nu}$ the energy-momentum tensor. The Landau and $\beta$-frame velocity are related, to leading order in $\Omega$, through $u^{\mu}_{L}=u^{\mu}+\sigma^{\omega}_{\varepsilon}\omega^{\mu}/(E+P)$. Applying a Lorentz boost along the $z$ axis brings the Landau frame velocity to the form corresponding to rigid rotation, such that the energy-momentum tensor and the charge currents read, to leading order in $\Omega$, as follows:
@@
The new Landau-frame vortical conductivities are related to the ones in the $\beta$ frame via
@@
where the last term represents the contribution from $\sigma^{\omega}_{\varepsilon}=Q_{A}$.
@@ 2.2 Non-conserved charges
Our work is concentrated on hydrodynamic systems that possess non-conserved charges associated with axial and helical degrees of freedom. The non-conservation of axial charge is caused by the axial anomaly in both Abelian (like QED) and non-Abelian (QCD) theories. In the massless quark limit, the anomaly leads to topologically induced dissipation in QCD due to either inter-vacuum tunnelling or thermal, sphaleron-induced processes related to the breaking of the axial symmetry by the axial anomaly itzykson80 ; bertlmann96 ; Kharzeev:2013jha ; buzzegoli20phd . The conservation of axial charge is also spoiled by an eventual non-zero fermionic mass, which, however, provides a relatively small effect for light quarks in heavy-ion collisions. In addition, the conservation of helical charge is also violated, both in QED and QCD, through the so-called helicity-violating pair annihilation (HVPA) processes (see Sec. 5.2 in Ref. Ambrus:2020oiw ).
@@
$\nabla_{\mu}J^{\mu}\equiv\frac{\partial Q}{\partial t}+{\bm{\nabla}}\cdot{\bm{J}}=-\frac{Q}{\tau}\,,$ (5)
where the relaxation rate is given by the appropriate relaxation time scale $\tau$. Equation (5) breaks Lorentz symmetry, thus indicating that the charge relaxation occurs only in the presence of matter and/or in thermal background.
@@
$\chi\delta\mu\rightarrow\chi_{\ell\ell^{\prime}}\delta\mu_{\ell^{\prime}}+\chi_{\ell T}\delta T\;,$ (7)
@@
$\nabla_{\mu}J^{\mu}_{\ell}=-\frac{1}{\tau_{\ell}}\left(\sum_{\ell^{\prime}=V,A,H}\chi_{\ell\ell^{\prime}}\delta\mu_{\ell^{\prime}}+\chi_{\ell T}\delta T\right)\,,\qquad\text{(no sum over $\ell=V,A,H$)}\,.$ (8)
@@
The conservation of the vector charge is encoded in the infinite vector relaxation time $\tau_{V}=\infty$. The other two relaxation times, $\tau_{A}$ and $\tau_{H}$, are, generally, finite quantities. Notice that despite the vector charge being a conserved quantity, in this prescription, its fluctuations affect the conservation of the other two charges so that the sum over $\ell^{\prime}$ in the right-hand-side of Eq. (8) runs over all three charges in the $V$, $A$, $H$ triad. While the proposal in Eq. (8) is in principle viable, detailed analysis within our $V$, $A$, $H$ system indicates that it leads to instabilities appearing at the level of the linear modes governing hydrodynamic fluctuations (see Appendix A for a more detailed discussion).
@@
Our preferred approach is to work on the basis of the chemical potential, which implies that the dissipation of the charge is controlled not by the magnitude of the charge density in the right-hand side of Eq. (5) but rather by the magnitude of the corresponding chemical potential. One argument supporting this approach is based on the Chiral Magnetic Effect, giving rise to the Chiral Magnetic Wave. Indeed, the non-conservation of the axial charge appears due to the presence of the anomalous triangular diagram with the axial-vector-vector (AVV) vertex corners. The anomalous generation of the vector current (one V-corner) along the magnetic field (another V-corner) is given by the same diagram, whose remaining A-corner directly links the magnitude of the generated current with the axial chemical potential and not the axial charge density. In this sense, the densities serve as the auxiliary thermodynamic characteristics of the system. At the same time, the chemical potentials play a primary role in the grand-canonical description of a system, while charge densities are derived quantities. In addition, the fact that a chemical potential for non-conserved charges is not well-defined, in a thermodynamic sense, quantity suggests that a system with non-conserved charges should be equilibrated to a state where such chemical potential is absent. Taking these considerations into account, we propose that Eq. (5) should be generalized as follows:
$\nabla_{\mu}J^{\mu}_{\ell}=-\frac{\chi_{\ell}}{\tau_{\ell}}\delta\mu_{\ell}\,,\qquad\text{(no sum over $\ell=V,A,H$)}\,,$ (9)
which amounts to taking only the diagonal element in Eq. (8).
@@
In summary, while the two approaches are identical in the case of a single-charge system [c.f. Eqs. (5) and (6)], they differ for systems that contain more than one charge. Since the charge dissipation is taken into account within an effective charge-relaxation-time formalism, we are free to choose between any of these prescriptions to proceed further. To this end, we notice that the first approach gives rise to unnatural unstable modes, as we point out in Appendix A. We will, therefore, focus our discussion only on the second approach.

### v2

@@ Abstract
In this paper, we consider the effect of interactions on the local, average polarization of a quantum plasma of massless fermion particles characterized by vector, axial, and helical quantum numbers. Due to the helical and axial vortical effects, perturbations in the vector charge in a rotating plasma can lead to chiral and helical charge transfer along the direction of the vorticity vector. At the same time, interactions between the plasma constituents lead to the dissipation of the helical charge through helicity-violating pair annihilation (HVPA) processes and of the axial charge through the axial anomaly. We will discuss separately a QED-like plasma, in which we ignore background electromagnetic fields and thus the axial charge is approximately conserved, as well as a QCD-like plasma, where instanton effects lead to the violation of the axial charge conservation, even in the absence of background chromomagnetic fields. The non-conservation of helicity and chirality leads to a gapping of the Helical, Axial, and mixed Axial-Helical vortical waves that prevents their infrared modes from propagating. On the other hand, usual dissipative effects, such as charge diffusion, lead to significant damping of ultraviolet modes. We end this paper with a discussion of the regimes where these vortical waves may propagate.
keywords:
Helicity, Vorticity, Dirac fermions, Vortical effects
@@ 1 Introduction
Vortical backgrounds – represented, for example, by whirlpools in fluids or plasmas – host a similar class of hydrodynamic phenomenon known under the common name of Chiral Vortical Effects Vilenkin:1978is ; Vilenkin:1979ui . These vortical transport effects generate intertwined oscillations of vector and axial currents, as well as their charges, propagating along the rotation axis Jiang:2015cva ; Gorbar:2017toh . The resulting Chiral Vortical Waves, which are supported by the presence of the chiral anomaly and its mixed chiral-gravitational counterpart Landsteiner:2011cp , were suggested to exist in the rotating quark-gluon plasma created in highly-vortical noncentral heavy-ion collisions Jiang:2015cva .
@@
The picture described above traditionally represents a chiral fluid as a two-component system that includes only vector and axial degrees of freedom that propagate coherently. However, at this point, it is worth noticing that fermions possess a third type of quantum number associated with the fermion helicity, which is distinct from the vector and axial charges Pal:2010ih . Helicity and chirality are distinguishable physical properties of fermions that are often inaccurately identified with each other Ambrus:2019ayb ; Ambrus:2019khr . The presence of this additional third characteristic of the fermion ensemble leads to the associated helical vortical transport phenomena that affect, in particular, the spectrum of the hydrodynamic modes Ambrus:2019khr . [^1] Here, for completeness, we mention that similar, helicity-related nontrivial transport and associated hydrodynamic modes in the magnetic field background has been uncovered in Refs. Ambrus:2023erf . In this paper, we concentrate only on rotating systems. Before proceeding further, it is worth mentioning that rotating media may also produce yet another, exotic “zilch” currents that are different from vector, chiral and helical charges. While the conserved zilch currents appear naturally in photonic gases Chernodub:2018era ; Prokhorov:2020npf , they can also emerge in system of chiral fermions Alexandrov:2020zsj . Below, we concentrate on the vector-axial-helical triad of quantum numbers, leaving the zilch charge aside.
@@
In our companion paper Morales-Tejera:2024uzg , the role of helical degrees of freedom in vortical chiral fluids is explored in an academic dissipationless limit when both axial and helical quantum numbers are conserved. This study has confirmed the presence of the Helical Vortical Wave Ambrus:2019khr , which corresponds to a gapless hydrodynamic excitation acting mainly in the helical and vector sectors of charge densities. Surprisingly, it also has uncovered yet another gapless wave in the system: the Axial Vortical Wave, which is driven by an axial imbalance accompanied, in the presence of vector (baryon) or helical (spin-polarized) background, by coherent fluctuations in axial and helical/vector charges and their currents. At high vector (baryon) densities, where the fluid becomes degenerate, both the Helical and the Axial Vortical Waves merge into a hydrodynamic excitation, the Axial-Helical Vortical Wave. The fascinating feature of the Axial Vortical Wave is its spatial non-reciprocity, which makes this hydrodynamic mode distinguishable from the usual acoustic waves: a pure Axial Vortical Wave propagates only opposite (along) the direction of the angular velocity for positive (negative) axial imbalance while being unable to travel in the opposite direction.
This paper aims to study the fate of the hydrodynamic vortical excitations in realistic plasmas, which are characterized by axial and helical charge relaxation as well as kinetic damping effects. A similar analysis has been extensively applied to determine possible experimentally observable signatures Ke:2012qb ; STAR:2015wza of a similar excitation, the Chiral Magnetic Wave, that emerges in the magnetic field background and incorporates oscillations of the vector and axial charges Newman:2005hd ; Kharzeev:2010gd .
@@
In addition to the standard axial charge relaxation effects – related to the fact that the axial charge is not a conserved quantity in quantum field theory – the vector sector is also subjected to relaxation effects caused by the electromagnetic backreaction Kharzeev:2010gd . The latter phenomenon is supported by a finite electrical conductivity of the plasma that leads to an efficient screening of the local vector charge fluctuations carrying nonvanishing electric charge density. As a result, the Chiral Magnetic Wave is essentially “overdamped” in realistic environments such as quark-gluon plasma: the relaxation length of the wave appears to be shorter than the half-period of the wave oscillation Shovkovy:2018tks . [^2] A recent work Ahn:2024ozz highlights the regimes, in which the Chiral Magnetic Wave can still be observed in a condensed matter setup as an underdamped collective mode in a Weyl semimetal which hosts relativistic fermionic quasiparticles similarly to quark-gluon plasma. In the vector-chiral sector of rotating plasma, a related analysis has been performed in Ref. Rybalka:2018uzh , where various waves were analyzed, including the longitudinal sound wave, a circularly polarized vortical wave and diffusive modes that involve an oscillating vector charge density.
@@
Here, we include in the analysis the helical charge, which enriches the spectrum of the hydrodynamic excitations. In particular, the Helical Vortical Wave Ambrus:2019khr involves the oscillations of the helical and vector charge densities, and, therefore, our analysis gives a higher, conservative bound of the wave propagation length caused by the helical charge relaxation and kinetic dissipation effects. The electromagnetic backreaction of the vector charge density, which is not considered in this paper, will restrict this length even further. On the other hand, the Axial and Axial-Helical Vortical waves, found in the companion paper Morales-Tejera:2024uzg , are driven essentially by electrically neutral helical and axial charge densities. Therefore, they are not subjected to the backreaction coming from the electromagnetic sector associated with the vector charge fluctuations.
@@
One may argue that the helical degree of freedom is as good (bad) as the chiral charge of massive fermions: neither of them is conserved in the physically relevant theories, such as massive QED or QCD. Chirality is not conserved for massless fermions either, due to instanton-like effects that are present in the QCD medium, even in the absence of background chromomagnetic fields. In the relaxation time approximation, the pertinence of these quantum numbers for the dynamics of the system is determined by the relative order of magnitudes of the axial, $\tau_{A}$, and helical, $\tau_{H}$, relaxation times, which show how fast these charges dissolve in the system. Similarly to the axial (chiral) charge Ruggieri:2016asg ; Astrakhantsev:2019zkr , the relaxation of the helical charge might also be a rather fast process Ambrus:2020oiw . In addition to the charge relaxation times, the system is also characterized by the kinetic relaxation time $\tau_{R}$, which encodes how fast fluctuations in the thermodynamic characteristics of the system (such as pressure and energy density) relax towards the thermodynamic equilibrium. In general, all three relaxation times, $\tau_{R}$, $\tau_{A}$, and $\tau_{H}$, are independent of each other. An expression for $\tau_{H}$ was derived in Ref. Ambrus:2019khr for a neutral plasma. In this paper, we will extend this derivation for the case of a finite vector imbalance, encoding a difference between the number of particles and anti-particles in the system, which is necessary in order to assess its relevance in the limit of a degenerate (dense) system. A similar calculation for the axial relaxation time is beyond the scope of this work; as such, an analysis must be performed in fully non-perturbative QCD using, e.g., lattice methods.
@@
The structure of the paper is as follows. In Sec. 2, we briefly present the setup for the study of linear perturbations in the quantum plasma with vector/axial/helical degrees of freedom, following the approach in Ref. Morales-Tejera:2024uzg , which we extend to the case when the helical and/or axial charges are not conserved. Our focus is on the case of an unpolarized background plasma, in which both the helical and axial chemical potentials vanish.
In Sec. 3, we take into account the non-conservation of the helical charge stemming from local processes using a relaxation time approximation. Such a scenario is relevant for a massless QED plasma, where both vector and axial charges are conserved due to an unbroken U(1) gauge symmetry and the vector nature of inter-particle interactions, respectively. The Chiral Vortical Wave, as considered in Gorbar:2017toh , is then recovered in the limit where the helical charge dissipates instantaneously. An interplay between Chiral and Helical Vortical Waves, as well as a distinct role played by the Axial Vortical Wave, are discussed.
@@ 2 Hydrodynamic description of waves in V/A/H fluids
In this section, we briefly review the anomalous transport in the presence of rotation. In Subsec. 2.1, we review the constitutive equations for the energy-momentum tensor and the charge currents for a Dirac fluid with vector, axial, and helical imbalance undergoing rigid rotation, as seen in the Landau (or energy) frame. Incorporating charge non-conservation is discussed in Subsec. 2.2. We then summarize the treatment of small perturbations around this rigidly rotating state in Subsec. 2.3.
@@ 2.1 Landau frame decomposition for slow rotation
$T^{\mu\nu}\simeq E_{\beta}u^{\mu}u^{\nu}-P_{\beta}\Delta^{\mu\nu}+u^{\mu}W_{\beta}^{\nu}+u^{\nu}W_{\beta}^{\mu},$
$J^{\mu}_{\ell}\simeq Q_{\ell;\beta}u^{\mu}+V_{\beta}^{\mu},$ (1)
where $W_{\beta}^{\mu}=\sigma^{\omega}_{\varepsilon;\beta}\omega^{\mu}$ and $V^{\mu}=\sigma^{\omega}_{\ell;\beta}\omega^{\mu}$ represent the heat flux and diffusion currents in the fluid rest frame, $\omega^{\mu}=\frac{1}{2}\varepsilon^{\mu\alpha\beta\gamma}u_{\alpha}\partial_{\beta}u_{\gamma}\simeq\Omega\delta^{\mu}{}_{z}$ is the vorticity four-vector and we neglected terms of quadratic or higher order in $\Omega$. In the above expression, $u^{\mu}\partial_{\mu}=\Gamma(\partial_{t}+\Omega\partial_{\varphi})$ is the four-velocity of a rigidly-rotating fluid, which defines the so-called $\beta$ (thermometer) hydrodynamic frame. The energy density $E_{\beta}$ is related to the pressure $P_{\beta}$ through the ultrarelativistic equation of state, $E_{\beta}=3P_{\beta}$. The charge densities $Q_{\ell;\beta}$ and the heat $\sigma^{\omega}_{\varepsilon;\beta}$ and charge $\sigma^{\omega}_{\ell;\beta}$ vortical conductivities, measured in the $\beta$ frame, can be obtained from the thermodynamic pressure $P_{\beta}$ via
$Q_{\ell;\beta}=\frac{\partial P_{\beta}}{\partial\mu_{\ell}},\ \ \ \sigma^{\omega}_{\varepsilon;\beta}=Q_{A;\beta},\ \ \ \sigma^{\omega}_{\ell;\beta}=\frac{1}{2}\frac{\partial Q_{\ell;\beta}}{\partial\mu_{A}}.$ (2)
To linear order in $\Omega$, one can employ the Landau hydrodynamic frame, by which the four-velocity $u_{L}^{\mu}$ satisfies $T^{\mu}{}_{\nu}u_{L}^{\nu}=Eu_{L}^{\mu}$, with $E$ being the energy density and $T^{\mu\nu}$ the energy-momentum tensor. The Landau and $\beta$-frame velocities are related, to leading order in $\Omega$, through $u^{\mu}_{L}=u^{\mu}+\sigma^{\omega}_{\varepsilon}\omega^{\mu}/(E+P)$. Applying a Lorentz boost along the $z$ axis brings the Landau frame velocity to the form corresponding to rigid rotation, such that the energy-momentum tensor and the charge currents read, to leading order in $\Omega$, as follows:
@@
The new Landau-frame vortical conductivities appearing above are related to the ones in the $\beta$ frame via
@@
where the last term represents the contribution from the $\beta$-frame heat conductivity, $\sigma^{\omega}_{\varepsilon}=Q_{A}$.
@@ 2.2 Non-conserved charges
Our work is concentrated on hydrodynamic systems that possess non-conserved charges associated with axial and helical degrees of freedom. The non-conservation of axial charge is caused by the axial anomaly in both Abelian (like QED) and non-Abelian (QCD) theories. In the massless quark limit, the anomaly leads to topologically induced dissipation in QCD due to either inter-vacuum tunneling or thermal, sphaleron-induced processes related to the breaking of the axial symmetry by the axial anomaly itzykson80 ; bertlmann96 ; Kharzeev:2013jha ; buzzegoli20phd . The conservation of axial charge is also spoiled by an eventual non-zero fermionic mass, which, however, provides a relatively small effect for light quarks in heavy-ion collisions. In addition, the conservation of helical charge is also violated, both in QED and QCD, through the so-called helicity-violating pair annihilation (HVPA) processes (see Sec. 5.2 in Ref. Ambrus:2020oiw ).
@@
$\partial_{\mu}J^{\mu}\equiv\frac{\partial J^{0}}{\partial t}+{\bm{\nabla}}\cdot{\bm{J}}=-\frac{Q}{\tau}\,,$ (5)
where the relaxation rate is given by the appropriate relaxation time scale $\tau$. In order to preserve Lorentz symmetry, the charge on the right hand side of Eq. (5) must be obtained via contraction with the fluid four-velocity, $Q=u_{\mu}J^{\mu}$, thus indicating that the charge relaxation occurs only in the presence of matter and/or in thermal background.
@@
$\delta Q\rightarrow\delta Q_{\ell}=\chi_{\ell\ell^{\prime}}\delta\mu_{\ell^{\prime}}+\chi_{\ell T}\delta T\;,$ (7)
@@
$\partial_{\mu}J^{\mu}_{\ell}=-\frac{1}{\tau_{\ell}}\left(\sum_{\ell^{\prime}=V,A,H}\chi_{\ell\ell^{\prime}}\delta\mu_{\ell^{\prime}}+\chi_{\ell T}\delta T\right)\,,$
[…]
@@
The conservation of the vector charge is encoded in the infinite vector relaxation time $\tau_{V}=\infty$. The other two relaxation times, $\tau_{A}$ and $\tau_{H}$ are, generally, finite quantities. Notice that despite the vector charge being a conserved quantity, in this prescription, its fluctuations affect the conservation of the other two charges so that the sum over $\ell^{\prime}$ in the right-hand-side of Eq. (2.2) runs over all three charges in the $V$, $A$, $H$ triad. While the proposal in Eq. (2.2) is in principle viable, detailed analysis within our $V$, $A$, $H$ system indicates that it leads to instabilities appearing at the level of the linear modes governing hydrodynamic fluctuations (see Appendix A for a more detailed discussion).
@@
Our preferred approach is to work on the basis of the chemical potential, which implies that the dissipation of the charge is controlled not by the magnitude of the charge density on the right-hand side of Eq. (5) but rather by the magnitude of the corresponding chemical potential. One argument supporting this approach is based on the Chiral Magnetic Effect, giving rise to the Chiral Magnetic Wave. Indeed, the non-conservation of the axial charge appears due to the presence of the anomalous triangular diagram with the axial-vector-vector (AVV) vertex corners. The anomalous generation of the vector current (one V-corner) along the magnetic field (another V-corner) is given by the same diagram, whose remaining A-corner directly links the magnitude of the generated current with the axial chemical potential and not the axial charge density. In this sense, the charge densities serve as the auxiliary thermodynamic characteristics of the system. At the same time, the chemical potentials play a primary role in the grand-canonical description of a system, while charge densities are derived quantities. In addition, the fact that a chemical potential for non-conserved charges is not a well-defined quantity, in a thermodynamic sense, suggests that a system with non-conserved charges should be equilibrated to a state where such chemical potential is absent. Taking these considerations into account, we propose that Eq. (5) should be generalized as follows:
$\partial_{\mu}J^{\mu}_{\ell}=-\frac{\chi_{\ell}}{\tau_{\ell}}\delta\mu_{\ell}\,,\qquad\text{(no sum over $\ell=V,A,H$)}\,,$ (9)
which amounts to taking only the diagonal element in Eq. (2.2). Note that the relaxation time approximation used to describe the non-conservation of charges is semi-phenomenological and a priori disconnected from the relaxation time approximation of Anderson and Witting Anderson:1974nyl in kinetic theory used to describe dissipative effects in Sec. 5.1. One may employ, e.g., the Shakhov extension of the RTA Ambrus:2024qsa to allow for independent relaxation times, however we do not pursue this avenue in the present work.
@@
In summary, while the two approaches are identical in the case of a single-charge system [c.f. Eqs. (5) and (6)], they differ for systems that contain more than one charge. Since the charge dissipation is taken into account within an effective charge-relaxation-time formalism, we are free to choose between any of these prescriptions to proceed further. To this end, we notice that the first approach gives rise to unnatural unstable modes, as we point out in Appendix A. We will, therefore, focus our discussion only on the second approach, summarized in Eq.(9) above.